# Robustness Optimization for Compact Free-electron Laser Driven by Laser Wakefield Accelerators

### v2

@@ Acknowledgements.
Appendix A Measurement of shock front
@@
The shock front is essential for the injection process in LWFA and thus significantly impacts the beam quality. As previously discussed, the position jitter with an RMS value of 4.9 μm was used in our simulations, which was experimentally measured. To capture the shock front properties, a probe laser passed through the plasma region and imaged using a 4f imaging system before entering the interferometer. The resulting interference patterns were captured using a charged-coupled device (CCD), with the shock front region depicted in Fig. 7(a). The corresponding phase distribution and the first-order derivative of the fringe phase are illustrated in Fig. 7(b) and Fig. 7(c), respectively. To characterize the position jitter of the shock front, we performed continuous acquisition of 100 experimental shots, as shown in Fig. 7(d). An RMS value of $\sigma_{zshock}=4.9$ μm was obtained. The statistical parameter is represented in the red shaded region in Fig. 7(d), providing quantitative analysis of the shock wave stability.

### v4

@@ Acknowledgements.
Appendix A Appendix A: Measurement of shock front
The shock front is essential for the injection process in LWFA and thus significantly impacts the beam quality. As previously discussed, the position jitter with an RMS value of 4.9 μm was used in our simulations, which was experimentally measured. To capture the shock front properties, a probe laser passed through the plasma region and imaged using a 4f imaging system before entering the interferometer. The resulting interference patterns were captured using a charged-coupled device (CCD), with the shock front region depicted in Fig. 6(a). The corresponding phase distribution and the first-order derivative of the fringe phase are illustrated in Figs. 6(b) and  6(c), respectively. To characterize the position jitter of the shock front, we performed continuous acquisition of 100 experimental shots, as shown in Fig. 6(d). An RMS value of $\sigma_{zshock}=4.9$ μm was obtained. The statistical parameter is represented in the red shaded region in Fig. 6(d), providing quantitative analysis of the shock wave stability.
@@ Appendix A Appendix A: Measurement of shock front
Appendix B Appendix B: Gain length optimization
$+51\eta_{d}^{0.95}\eta_{\gamma}^{3}+5.4\eta_{d}^{0.7}\eta_{\epsilon}^{1.9}+1140\eta_{d}^{2.2}\eta_{\epsilon}^{2.9}\eta_{\gamma}^{3.2}\;$
$\eta_{d}$ $=$ $\frac{L_{G0}}{2k_{r}\sigma_{x}\sigma_{y}}\;,$ (3)
$\eta_{\gamma}$ $=$ $\frac{L_{G0}}{4\pi\lambda_{u}\sigma_{\eta}}\;.$ (5)
In the optimization framework targeting $<L_{G}^{-1}>$, we conducted parallel particle tracking simulations utilizing the 13-electron-beam ensemble illustrated in Fig. 1. The resulting transverse beam characteristics were subsequently employed to calculate $<L_{G}^{-1}>$, which serves as the primary quantitative metric for objective function evaluation. The parameter ranges for the optimization are presented in Table 2.
Table 2: Parameter space used in $<L_{G}^{-1}>$
optimization. Parameter Min. value Max. value Best value [^2] Note $l_{0}-l_{6}$ in units of meters (m), $q_{1}-q_{6}$ in units of per square meter (m-2). $l_{0}$ 0.03 1.0 0.08 $l_{1}$ 0.03 1.0 0.11 $l_{2}$ 0.03 1.0 0.16 $l_{3}$ 0.3 3.0 1.17 $l_{4}$ 0.3 3.0 2.98 $l_{5}$ 0.3 3.0 0.39 $l_{6}$ 0.3 3.0 0.33 $q_{1}$143 143 -139 $q_{2}$143 143 141 $q_{3}$143 143 -69 $q_{4}$45 45 -0.34 $q_{5}$45 45 10.51 $q_{6}$45 45 -10.82
Figure 7: The optimization results of the $<L_{G}^{-1}>$. (a) The simulated objective function $<L_{G}^{-1}>$ (dots) with the cumulative best results (black dashed line). (b)-(c) The evolution of the beam size (b) and the normalized projected emittance (c) along the beamline in horizontal (blue) and vertical (orange) directions. The tracking beamline was configured using the best parameters obtained from the optimization.
Figure 7(a) illustrates the evolution of the objective function $<L_{G}^{-1}>$ throughout the BO process, with the cumulative optimal results represented by the black dashed line. The optimization commenced with random sampling within the predefined parameter space detailed in Table 2. A progressive convergence of the objective function emerges after approximately 1200 iterations, with the optimal solution achieved at iteration 10,071, corresponding to an average 3D power gain length of approximately 0.24 m. The resulting beam sizes and transverse emittance for the 13-electron-beam ensemble under the optimized conditions are presented in Figs. 7(b) and  7(c), respectively. It should be noted that slice-averaged energy spread of the e beam is employed when calculating the power gain length. The radiation gain predominantly originates from specific slices with advanced beam qualities, while other slices contribute minimally to the overall gain. The optimization based on slice-averaged values may cause potential deviations, considering the different properties over slices of the e beam from LWFAs. To address this limitation, the optimization using the radiation energy as an alternative objective function is implemented.